# What is glacier sliding?

### v1

@@ Abstract
Glacier and ice-sheet motion is fundamental to glaciology. However, there is still no clear consensus for the optimal way to describe glacier and ice-sheet sliding. Typically, sliding is parameterised using a traction coefficient nominally linked to a given theory describing one or a limited set of sliding processes. However, this approach precludes the possibility of multiple simultaneous and spatio-temporally varying sliding modes with inaccuracies resulting in model error propagation as the system evolves away from the conditions under which it was optimised for. Here, revisiting early theoretical work,we describe glacier sliding as a scale- and setting-dependent ‘inner flow’ that arises from multiple overlapping sub-processes, bridging divides between hard and soft beds, rough and smooth beds, and stick-slip and continuous sliding as well as providing a consistent definition for form drag. The corresponding ‘outer flow’ then accounts for ‘normal’ ice deformation. We propose that the significance of ‘Iken’s bound’ is then reduced if form drag dominates over subglacial cavitation in a given region, possibly explaining the persistent functionality of Weertman-type sliding in process-agnostic sliding studies over rough topography. Last, reviewing observation-based studies we suggest that a simple ‘unified’ sliding relationship controlled by a single tunable coefficient may not be a realistic prospect, but that a Weertman-type relationship with careful consideration of the power value is presently the ‘least bad’ option.
@@ 1 Background
Glacier sliding is a phenomenologically distinct problem within earth sciences, separated from simpler plane-on-plane sliding problems between two elastic solids by multiple confounding factors. The relative softness of ice allows for both viscous deformation under a non-Newtonian power law and discrete slip at the ice-bed interface, which may occur in a stick-slip or continuous manner. A ‘soft’ sediment phase provides a further medium to convey deformation in many instances (and is shaped in turn by glacial processes), while ‘hard’ underlying bedrock can constitute a highly topographically irregular base ranging from <1 cm asperities to dramatic fjords, necessitating ice deformation around rigid obstacles over length scales covering 5 orders of magnitude. A subglacial hydrological system comprised largely of melted ice may significantly alter the effective pressure of the ice upon its substrate, and influence the roughness and material characteristics of the base itself. Debris entrained within basal ice interacts with the subglacial environment and influences both basal ice properties and the frictional properties of the glacier sole. These considerations present a wide parameter space for basal boundary conditions, which may also exhibit considerable spatial and temporal variations. The net outcome is significant difficulty in the derivation of a ‘unified’ or ‘generalised’ sliding relationship, usually expressed in the form
@@
where $\tau_{b}$ is basal traction, $u_{b}$ is ice velocity tangential to the ice-bed interface, $N=P_{i}-P_{w}$ is effective pressure, $P_{w}$ is subglacial water pressure, and $P_{i}$ is the ice overburden pressure.
Determining an appropriate sliding relationship is central to producing tractable, physically-based projections of the contribution of ice sheets and glaciers to sea-level rise and fresh-water fluxes over the coming centuries. Extant sliding relationships (Fig. 1, Table. 1) influence ice-sheet model output at catchment to entire ice sheet scales, with the distinction between bounded traction (plastic, regularised-Coulomb) and unbounded traction (all other relationships) being particularly consequential (Gillet-Chaulet and others,, 2012; Parizek and others,, 2013; Ritz and others,, 2015; Tsai and others,, 2015; Bons and others,, 2018; Kyrke-Smith and others,, 2018; Brondex and others,, 2019; Åkesson and others,, 2021; Lippert and others,, 2024; Trevers and others,, 2024), yet around five [^1] Some sliding relationships can be considered a simplified versions of one another. For example Weertman-type and pseudo-plastic can be manipulated into linear or plastic relationships, the $N$ in Budd sliding is effectively subsumed into the $C$ of Weertman-type sliding during inversion if it is assumed that subglacial water pressure does not vary in time, making this number subjective. (Section 3) sliding relationships implicitly or explicitly capturing different sliding processes are used in the Ice Sheet Model Intercomparison Project (ISMIP) models (Goelzer and others,, 2020; Seroussi and others,, 2020), which guide worldwide policy on sea level rise mitigation. This state of affairs is not immediately straightforward to resolve, as the inverse methods that tune for $C$, the ‘traction coefficient’, will produce reasonable basal stress states whichever equation in Table 1 is chosen as the global stress balance condition must be met (Joughin and others,, 2009; Arthern and Gudmundsson,, 2010). There is therefore no simple approach to discounting sliding relationships given solely a snapshot of glacier or ice sheet velocity and geometry. Errors associated with an inappropriate choice of basal sliding relationship subsequently increase as a function of model run time (Aschwanden and others,, 2021).
Figure 1: Existing sliding parameterisations and how their development relates to soft- and hard-bed theory. Traction and velocity values are plausible, but for illustrative purposes only.
@@
Table 1: Selected sliding relationships applied in numerical glacier and ice-sheet models, expressed in one-dimensional form. Denotes that the sliding relationship is used in Goelzer and others, (2020) or Seroussi and others, (2020) ISMIP experiments. In all equations $\tau_{b}$ is basal traction and $u_{b}$ is the basal velocity tangential to the bed. Where present, $n$ is the exponent used in Glen’s flow law (usually 3) and is only used where an explicit link to Glen’s flow law is made in the paper proposing the sliding relationship, $m$ is an exponent often related to Glen’s flow law, but not always explicitly, $N$ is the effective pressure, $C$ is the traction coefficient which may be adapted from its use in the original paper for simplified intercomparison (for example we replace $1/C$ with $C$ in the Budd relationship), $\tau_{c}$ is the yield stress in the pseudo-plastic relationship, $q$ is a fitting parameter in the Budd relationship, and $u_{t}$ is the threshold velocity in the pseudo-plastic and regularised-Coulomb relationships. The rate-weakening 2D cavities relationship in Fig. 1 is not included, as no numerical ice-sheet or glacier models include rate-weakening behaviour above a given velocity threshold. Depending on convention, the RHS may be negative in the original paper to indicate a traction force opposite to velocity direction. In dimensions higher than 1 the sliding relationship is vectorised. Notes: Linear: Adapted from Nye, (1969) and Kamb, (1970). As applied in e.g. Morlighem and others, (2013). The version with $N$ is used in some Goelzer and others, (2020) experiments. Weertman: Equation featuring $n$ is adapted from Weertman, (1957) where $n$ is set as 3. $m$ is often used in models instead of $n$ where $m$ is usually between 3 and 4. Pseudo-plastic: Effectively the same as Weertman but included for frequent use in PISM applications e.g. Aschwanden and others, (2016), the pseudo-plastic relationship can be varied between plastic ($m=\infty$), linear ($m=1$), or Weertman-type ($2\leq m\leq 4$) behaviour. Regularised-Coulomb: The simplest form excluding $N$, adapted from (Joughin and others,, 2019). Other formulations exist such as $\tau_{b}=Cu_{b}N\left(\frac{{u_{b}}^{-n+1}}{u_{b}+A_{s}C^{n}{N}^{n}}\right)^{1/n}$ (Helanow and others,, 2021) where $A_{s}$ is a parameter specific to this formulation. The differences between regularised-Coulomb relationships are not negligible (Appendix A). Plastic: As used in e.g. Bougamont and others, (2014) and the PISM default (Winkelmann and others,, 2011). Budd: As developed in Budd and others, (1984) where $q=1$ and $m=3$ and used in Budd and Jacka, (1989) with $q=2$ and $m=1$. The alternative and simplified Budd formulation is as used in Choi and others, (2022). The proposal of Tsai and others, (2021) also bears similarities with Budd sliding.
@@
In numerical ice-sheet models, these sliding relationships form a critical thread between the processes occurring at the sub-grid-scale and the numerical solution defined at the grid scale (where ‘grid’ refers to discretised model grid cells or mesh elements)  [^2] This was not always the case, with purely analytical sliding relationships predating the first numerical models by decades (Weertman,, 1957; Rasmussen and Campbell,, 1973; Jenssen,, 1977) which may explain why sub-grid-scale processes were not initially a concern.. In discretised continuum mechanics models, the distinction between sub-grid-scale and grid-scale is usually made using a ‘representative volume element’ that is large enough to be representative of the bulk behaviour of micro-scale processes within the constitutive material it represents. For ice deformation, this is straightforward – an ice volume of 1 m3, much below the resolution of even the most detailed numerical models, will contain sufficient variability in crystal orientation, size, and impurities as to have effectively the same response to an applied stress field as a different 1 m3 of the same material. Lab-based experiments using small volumes of ice (e.g. Glen,, 1952) can therefore be reasonably applied to much larger-scale models, even if separate issues such as the importance of tertiary creep remain (Adams and others,, 2021).
@@
There is not, however, an obvious representative volume, or ‘representative surface’ element size for glacier sliding processes, with investigations into glacier sliding covering a disparate range of scales extending over seven orders of magnitude from millimetres to 10s of kilometres (Fig. 2). For example, lab experiments for ice-sediment shearing behaviour may be conducted at a scale of tens of centimetres (e.g. Iverson and others,, 1998; Zoet and Iverson,, 2020), up-scaled to field settings covering an entire glacier where subglacially entrained clasts, variable subglacial hydrology, or bedrock obstacles further influence the relationship between traction and velocity (e.g. Iverson and others,, 1995 Hedfors and others,, 2003, Gimbert and others,, 2021), and ultimately applied to ice-sheet models where individual grid cells can exceed 10s of kilometres (or have a highly variable size across a given domain) and cover an uncertain and/or heterogeneous subglacial landscape (e.g. Kyrke-Smith and others,, 2018; Holschuh and others,, 2020; Paxman and others,, 2021). Nonetheless, routinely-used sliding parameterisations in ice sheet models almost exclusively use theories derived for one or a limited set of sliding processes at small scales (Goelzer and others,, 2020; Seroussi and others,, 2020), without representing the full range of scale-dependent processes.
@@
Figure 2: Spatial scales for theories and parameterisations of basal sliding. A solid line indicates coverage as explicitly defined in the associated paper paper while a dashed line indicates probable situation dependent coverage and an arrow indicates extension beyond the scale bar or an uncertain coverage beyond the given limit. Reasoning behind the positioning of each spatial range is provided in Appendix B.
@@
In this paper we provide a clear framework for describing glacier sliding as behaviour emerging from multiple scale- and setting-dependent processes (summarised in Fig. 4) rather than as a single dominant process. We reconsider the ‘inner-outer’ flow division (Section 2) first proposed by Fowler, (1977), but with a more universal approach to the inclusion of numerous sliding sub-processes (Section 3). We briefly cover standard soft- and hard-bed theory (Sections 3.1 and 3.2); form drag, roughness, and temperate ice (Section 3.3); stick-slip vs. continuous sliding (Section 3.4); the overlap between soft- and hard-bed sliding (Section 3.5); the basal ice layer and ice rheology (Section 3.6); and how these sub-processes are modified by subglacial hydrology (Section 4). We focus on annual timescales as these are the most relevant for predictive and palaeo models, and largely avoid the transient influence of e.g. supraglacial lake drainages (Das and others,, 2008), rainfall events (Doyle and others,, 2015), or seasonal modulation (Bartholomew and others,, 2011; Sole and others,, 2013). In Section 5 we review process-agnostic studies fitting sliding relationships to observational data. Last, we consider if a truly unified sliding relationship is a realistic possibility, consider the settings under which Iken’s bound can be considered to hold, and make suggestions for the application of sliding relationships to models (Section 6).
@@
Figure 4: Schematics of glacier sliding sub-processes and controls discussed in the text. Scale is intentionally omitted but scale generally increases left to right and top to bottom through sub processes. Hydrology in k and l is considered as a process modifier, as no ice mass is transported and these are distinguished with blue outline boxes. a Form drag over sedimentary clasts. b Shear of sediment. c Ice with clasts sliding over a flat hard bed. d Stick-slip events. e Regelation. f Deformation within the basal boundary layer. g Cavitation. h Spatially variable sliding rates over rough topography. i Spatially variable ice deformation over rough topography. j Temperate layer processes. k Hydrology and channelisation. l Subglacial lakes.
@@ 2 Inner-outer flows
In our application, the inner flow accounts for the relevant subset of processes outlined in Fig. 4 and Section 3 (as opposed to the simpler set of processes considered in Fowler1977) and the outer flow comprises englacial ice deformation with minimal influence from the processes in Fig. 4. In a modelling sense, the outer flow represents all motion not captured by the basal sliding component. A more complex example of multiple sub-processes operating in an inner flow is given in Appendix C.
@@
A distinction between an inner and outer flow begins to form an inclusive description of glacier sliding, but a consistent definition for the position of the inner-outer boundary does not immediately materialise – even if an explicit inner-outer boundary is in keeping with its implicit use in previous studies (e.g. Schoof and Clarke,, 2008; Zoet and Iverson,, 2020; Helanow and others,, 2021). The inner flow does not neatly fit the idea of a boundary layer as typically used in fluid dynamics to describe a region of turbulent flow separation (Batchelor,, 1967 and discussed further in Section 3.3) and shares only broad similarities with a bottom boundary layer as used in oceanography models (Campin and Goosse,, 1999). A numerical approach also means that the inner and outer flow can not be matched analytically as they are in Fowler1977. However, perturbations in strain rates tend to decrease in amplitude with increasing height above the bed, either exponentially (Balise and Raymond,, 1985) or as a power law (Chandler and others,, 2006); corresponding length scales for the decay of inner layer strain rate perturbations could then form a basis for defining the inner-outer layer boundary.

### v3

@@ Abstract
Glacier and ice-sheet motion is fundamental to glaciology. However, there is still no clear consensus for the optimal way to describe the sliding component of glacier and ice-sheet dynamics. Typically, sliding is parametrised using a traction coefficient linked to a given theory describing one or a limited set of sliding processes. However, this precludes the possibility of multiple overlapping and spatio-temporally varying sliding modes with inaccuracies resulting in model error propagation as the system evolves away from the conditions under which it was optimised for. Here, we describe glacier sliding as a scale- and setting-dependent ’inner flow’ that arises from multiple overlapping sub-processes, including viscous ice deformation and obstacle resistance (or ’form drag’). The corresponding ’outer flow’ then accounts for ice deformation that is only minimally influenced by bed properties. Following this framework, we suggest that a rough-smooth bed division may be more consequential than a hard-soft one, and that the importance of ’Iken’s bound’ can be significantly reduced if viscous ice deformation becomes a significant process in a given region within the inner flow, which may explain the persistent functionality of power-law sliding in large-scale process-agnostic sliding studies over rough topography. Further, reviewing observation-based studies and considering the diversity of sliding processes and settings, we suggest that a simple ’unified’ sliding relationship controlled by a single tunable coefficient may not be realistic. However, we show that given reasonable assumptions, sliding relationships should generally fall within a sum of regularised-Coulomb and power-law components, and suggest that a or compound, or power-law relationship with careful consideration of the power value may provide the most flexible way to account for the varied net effect of compound sliding sub-processes.
@@ 1 Background
Glacier sliding is a phenomenologically distinct problem within earth sciences, separated from simpler plane-on-plane sliding problems between two elastic solids by multiple confounding factors. The relative softness of ice allows for both viscous deformation under a non-Newtonian power-law and discrete slip at the ice-bed interface, which may occur in a stick-slip or continuous manner. A ‘soft’ sediment phase provides a further medium to convey deformation in many instances (and is shaped in turn by glacial processes), while ‘hard’ underlying bedrock can constitute a highly topographically irregular base ranging from <1 cm asperities to dramatic fjords, necessitating ice deformation around rigid obstacles over length scales covering 5 orders of magnitude. A subglacial hydrological system comprised largely of melted ice may significantly alter the effective pressure of the ice upon its substrate, and influence the roughness and material characteristics of the base itself. Debris entrained within basal ice interacts with the subglacial environment and influences both basal ice properties and the frictional properties of the glacier sole. These considerations present a wide parameter space for basal boundary conditions, which may also exhibit considerable spatial and temporal variations. The net outcome is significant difficulty in the derivation of a ‘unified’ or ‘generalised’ sliding relationship, usually expressed in the form
@@
where $\tau_{b}$ is basal traction, $u_{b}$ is ice velocity tangential to the ice-bed interface, $N=P_{i}-P_{w}$ is effective pressure, $P_{w}$ is subglacial water pressure, and $P_{i}$ is the ice overburden pressure (or with $u_{b}$ and $\tau_{b}$ vectorised for higher dimensions, e.g. Eq. 9).
Determining an appropriate sliding relationship is central to producing tractable, physically-based projections of the contribution of ice sheets and glaciers to sea-level rise and fresh-water fluxes over the coming centuries. Extant sliding relationships (Fig. 1, Table. 1) influence ice-sheet model output at catchment to entire ice sheet scales, with the distinction between bounded traction (plastic, regularised-Coulomb) and unbounded traction (all other relationships) being particularly consequential (Gillet-Chaulet and others,, 2012; Parizek and others,, 2013; Ritz and others,, 2015; Tsai and others,, 2015; Bons and others,, 2018; Kyrke-Smith and others,, 2018; Lippert and others,, 2024; Trevers and others,, 2024), resulting in up to a 100% increase in mass loss over 100 years dependent on initialisation (Brondex and others,, 2019; Åkesson and others,, 2021). Nonetheless, around five [^1] Some sliding relationships can be considered a simplified versions of one another. For example power-law and pseudo-plastic can be manipulated into linear or plastic relationships, the $N$ in Budd sliding is effectively subsumed into the $C$ of power-law sliding during inversion if it is assumed that subglacial water pressure does not vary in time, making this number subjective. (Section 3) sliding relationships implicitly or explicitly capturing different sliding processes are used in the Ice Sheet Model Intercomparison Project (ISMIP) models (Goelzer and others,, 2020; Seroussi and others,, 2020), which guide worldwide policy on sea level rise mitigation. This state of affairs is not immediately straightforward to resolve, as the inverse methods that tune for $C$, the ‘traction coefficient’, will produce reasonable basal stress states whichever equation in Table 1 is chosen as the global stress balance condition must be met (Joughin and others,, 2009; Arthern and Gudmundsson,, 2010). There is therefore no simple approach to discounting sliding relationships given solely a snapshot of glacier or ice sheet velocity and geometry. Errors associated with an inappropriate choice of basal sliding relationship subsequently increase as a function of model run time (Aschwanden and others,, 2021).
Figure 1: Existing sliding parameterisations and how their development relates to soft- and hard-bed theory. R-c+P-l represents Regularised-Coulomb plus power-law. Traction and velocity values are plausible, but for illustrative purposes only.
@@
Table 1: Selected sliding relationships applied in numerical glacier and ice-sheet models, expressed in one-dimensional form. Denotes that the sliding relationship is used in Goelzer and others, (2020) or Seroussi and others, (2020) ISMIP experiments. In all equations $\tau_{b}$ is basal traction and $u_{b}$ is the basal ice velocity tangential to the bed. Where present, $n$ is the exponent used in Glen’s flow law (usually 3) and is only used where an explicit link to Glen’s flow law is made in the paper proposing the sliding relationship, $m$ is an exponent often related to Glen’s flow law, but not always explicitly, $N$ is the effective pressure, $C$ is the traction coefficient which may be adapted from its use in the original paper for simplified intercomparison (for example we replace $1/C$ with $C$ in the Budd relationship), $\tau_{c}$ is the yield stress in the pseudo-plastic relationship, $q$ is a fitting parameter in the Budd relationship, and $u_{t}$ is the threshold velocity in the pseudo-plastic and regularised-Coulomb relationships. The rate-weakening 2D cavities relationship in Fig. 1 is not included, as no numerical ice-sheet or glacier models include rate-weakening behaviour above a given velocity threshold. Depending on convention, the RHS may be negative in the original paper to indicate a traction force opposite to velocity direction. In dimensions higher than 1 the sliding relationship is vectorised (e.g. Eq. LABEL:eq:slipvel). Notes: Linear: Adapted from Nye, (1969) and Kamb, (1970). As applied in e.g. Morlighem and others, (2013). The version with $N$ is used in some Goelzer and others, (2020) experiments. Weertman: Equation featuring $n$ is adapted from Weertman, (1957) where $n$ is set as 3. $m$ is often used in models instead of $n$ where $m$ is usually between 3 and 4. Pseudo-plastic: Effectively the same as Weertman but included for frequent use in PISM applications e.g. Aschwanden and others, (2016), the pseudo-plastic relationship can be varied between plastic ($m=\infty$), linear ($m=1$), or power-law ($2\leq m\leq 4$) behaviour. Regularised-Coulomb: The simplest form excluding $N$, adapted from (Joughin and others,, 2019). Other formulations exist such as $\tau_{b}=Cu_{b}N\left(\frac{{u_{b}}^{-n+1}}{u_{b}+A_{s}C^{n}{N}^{n}}\right)^{1/n}$ (Helanow and others,, 2021) where $A_{s}$ is a parameter specific to this formulation. The differences between regularised-Coulomb relationships are not negligible (Fig. A1). Plastic: As used in e.g. Bougamont and others, (2014) and the PISM default (Winkelmann and others,, 2011). Budd: As developed in Budd and others, (1984) where $q=1$ and $m=3$ and used in Budd and Jacka, (1989) with $q=2$ and $m=1$. The alternative and simplified Budd formulation is as used in Choi and others, (2022). Tsai and others, (2021) also produces a relationship similar to Budd sliding.
@@
In numerical ice-sheet models, these sliding relationships form a critical thread between the processes occurring at the sub-grid-scale and the numerical solution defined at the grid scale (where ‘grid’ refers to discretised model grid cells or mesh elements)  [^2] This was not always the case, with purely analytical sliding relationships predating the first numerical models by decades (Weertman,, 1957; Rasmussen and Campbell,, 1973; Jenssen,, 1977) which may explain why sub-grid-scale processes were not initially a concern.. In discretised continuum mechanics models, the distinction between sub-grid-scale and grid-scale is usually made using a ‘representative volume element’ that is large enough to be representative of the bulk behaviour of micro-scale processes within the constitutive material it represents. For ice deformation, this is straightforward – an ice volume of 1 m3, much below the resolution of even the most detailed numerical models, will contain sufficient variability in crystal orientation, size, and impurities as to have effectively the same response to an applied stress field as a nearby 1 m3 of the same material. Lab-based experiments using small volumes of ice (e.g. Glen,, 1952) can therefore be reasonably applied to much larger-scale models, even if separate issues such as the importance of tertiary creep remain (Adams and others,, 2021).
@@
There is not, however, an obvious representative volume, or ‘representative surface’ element size for glacier sliding processes, with investigations into glacier sliding covering a range of scales extending over seven orders of magnitude from millimetres to 10s of kilometres (Fig. 2). For example, lab experiments for ice-sediment shearing behaviour may be conducted at a scale of tens of centimetres (e.g. Iverson and others,, 1998; Zoet and Iverson,, 2020), up-scaled to field settings covering an entire glacier where subglacially entrained clasts, variable subglacial hydrology, or bedrock obstacles further influence the relationship between traction and velocity (e.g. Iverson and others,, 1995 Hedfors and others,, 2003, Gimbert and others,, 2021), and ultimately applied to ice-sheet models where individual grid cells can exceed 10s of kilometres (or have a highly variable size across a given domain) and cover an uncertain and/or heterogeneous subglacial landscape (e.g. Kyrke-Smith and others,, 2018; Holschuh and others,, 2020; Paxman and others,, 2021). Nonetheless, routinely-used sliding parameterisations in ice sheet models almost exclusively use theories derived for one or a limited set of sliding processes at small scales (Goelzer and others,, 2020; Seroussi and others,, 2020), without representing the full range of scale-dependent processes.
@@
Figure 2: Spatial scales for theories and parameterisations of basal sliding. A solid line indicates coverage as explicitly defined in the associated paper paper while a dashed line indicates probable situation dependent coverage and an arrow indicates extension beyond the scale bar or an uncertain coverage beyond the given limit. Reasoning behind the positioning of each spatial range is provided in Appendix A1.
@@
In this paper we provide a clear framework for describing glacier sliding as behaviour emerging from multiple scale- and setting-dependent processes (summarised in Fig. 4), rather than as a single dominant process. We reconsider the ‘inner-outer’ flow division (Section 2) first proposed by Fowler, (1977), but with a more universal approach to the inclusion of numerous sliding sub-processes over three dimensions and resistance from sub-grid-scale topography (Section 3). We briefly cover standard soft- and hard-bed theory (Sections 3.1 and 3.2); form drag, roughness, and temperate ice (Section 3.3); stick-slip vs. continuous sliding (Section 3.5); the overlap between soft- and hard-bed sliding (Section 3.6); the basal ice layer and ice rheology (Section 3.4); and how these sub-processes are modified by subglacial hydrology (Section 4). We focus on annual timescales as these are the most relevant for predictive and palaeo models, and largely avoid the transient influence of e.g. supraglacial lake drainages (Das and others,, 2008), rainfall events (Doyle and others,, 2015), or seasonal modulation (Bartholomew and others,, 2011; Sole and others,, 2013). In Section 5 we review process-agnostic studies fitting sliding relationships to observational data. Last, we consider if a simple (i.e. controlled by a single parameter) unified sliding relationship is a realistic possibility, consider the settings and scales under which Iken’s bound may not hold, and make suggestions for the application of sliding relationships to models (Section 6).
@@
Figure 4: Schematics of glacier sliding sub-processes and controls discussed in the text. Scale is intentionally omitted but scale generally increases left to right and top to bottom through sub-processes. Hydrological processes in k and l are considered as controls, since they do not contribute directly to ice transport; these are distinguished with blue outline boxes. a Form drag over sedimentary clasts. b Shear of sediment. c Ice with clasts sliding over a flat hard bed. d Stick-slip events. e Regelation. f Deformation within the basal boundary layer. g Cavitation. h Spatially variable sliding rates over rough topography. i Spatially variable ice deformation over rough topography. j Temperate layer processes. k Hydrology and channelisation. l Subglacial lakes.
@@ 2 Inner-outer flows
In our application, the inner flow accounts for the relevant subset of processes outlined in Fig. 4 and Section 3 (as opposed to the simpler set of processes considered in Fowler1977) and the outer flow comprises englacial ice deformation with minimal influence from the processes in Fig. 4. In a modelling sense, the outer flow represents all motion not captured by the basal sliding component. A more complex example of multiple sub-processes operating in an inner flow is given schematically in Fig. A2.
@@
An inner-outer flow distinction may facilitate a more holistic description of glacier sliding, but a consistent definition for the position of the inner-outer boundary does not immediately materialise – even if an explicit inner-outer boundary is in keeping with its implicit use in previous studies (e.g. Schoof and Clarke,, 2008; Zoet and Iverson,, 2020; Helanow and others,, 2021). The inner flow does not neatly fit the idea of a boundary layer as typically used in fluid dynamics to describe a region of turbulent flow separation (Batchelor,, 1967 and discussed further in Section 3.3) and shares only broad similarities with a bottom boundary layer as used in oceanography models (Campin and Goosse,, 1999). A numerical approach also means that the inner and outer flow can not be matched analytically as they are in Fowler1977. However, perturbations in strain rates tend to decrease in amplitude with increasing height above the bed, either exponentially (Balise and Raymond,, 1985) or as a power-law (Chandler and others,, 2006); corresponding length scales for the decay of inner layer strain rate perturbations could then form a basis for defining the inner-outer layer boundary.